How do you solve for y in the following equations:
-3/4x+y/3=5/12
x/5-y=4/5

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

(-3x/4)+y/3=5/12
(-9x+4y)/12= 5/12
multiply by 12
-9x+4y=5......................1
x/5-y=4/5
(x-5y)/5 = 4/5
multiply by 5
x-5y = 4......................2
..
multiply equation 2 by 9
9x-45y=36
add to (1)
-45y+4y=36+5
-41y=41
/-41
y =-1
..
plug value of y in 1
-9x-4=5
-9x= 4+5
/-9
x=-1
...
m.ananth@hotmail.ca

Answer by stanbon(75887)

-3/4x+y/3=5/12
x/5-y=4/5
---
Get rid of the fractions:
-9x+4y = 5
x - 5y = 4
--------
Solve the 2nd equation for "x":
x = 5y+4
Substitute and solve for "y":
-9(5y+4) + 4y = 5
-45y-36 +4y = 5
-------------
-41y = 41
y = -1
===================
Substitute into x = 5y+4 to solve for "x":
x = 5(-1)+4
x = -1
----------------
Cheers,
Stan H.